# Is uniform expressivity too restrictive? Towards efficient expressivity of GNNs

### v1

@@ 6.2 Experimental setup
Computing and approximating queries. For the first experiment, our set of input graphs is composed of pairs of trees $(T[0,k],T[1,k])_{k\in\mathbb{N}}$ presented in Section 4. We compare for $i\in[2]$, the two following GNNs:
  • i)
    The ReLU GNN that exactly computes the query $Q_{i}$.
    For $Q_{1}$, the ReLU GNN has sum-aggregation, $4$ iterations (depth of $Q$). The input and output dimensions are set for convenience to $4$ to have a constant dimension for all iterations. In particular, all features vectors of the input vertices are set to $e_{1}$, the first canonical vector of $\mathbb{R}^{4}$. The weights and biases are set accordingly to the construction in Section 5 (specified in Appendix C). Their exact values are explicited in the Appendix D.
    $\text{For every $t\in[4]$, and every $x,y\in\mathbb{R}^{4}$ }\quad\mathsf{comb}_{t}(x,y):=\operatorname{ReLU}(Ax+By+c)$
    For $Q_{2}$, the ReLU GNN has $7$ iterations.
  • ii)
    The Pfaffian GNN that expresses $Q_{i}$ non uniformly, with activation function: $\sigma:x\mapsto\frac{4}{\pi}\arctan(x)$, and sum-aggregation, with the same construction as in the proof of Theorem 5.1. Each combination function is given by a NN with $m$ layers, and one neuron per layer. The bias of the first layer is set to $-\frac{1}{2}$. In the last non-hidden layer (no activation), the weight is set to $\frac{1}{2}$, and the bias is set to $0.5$ to the last layer.
    Setting, $f_{m}:=\sigma^{m}(.)$, the neural network computes the function $g_{m}:=\frac{1}{2}+\frac{1}{2}f_{m}(.-1/2)$. Note that $g_{m}$ is indeed a function computed by a neural network with activation function $\sigma$ by adding a second neuron in each layer, setting the weights to 00, and biases to 00 except in the last layer where the bias is set to $1$ and the weight set to $\frac{\pi}{4\arctan(1)}$ to add a bias of $\frac{1}{2}$.
    $\text{For every $t\in[4]$, and every $x,y\in\mathbb{R}^{4}$ }\quad\mathsf{comb}_{t}(x,y):=g_{m}(Ax+By+c)$
    where $A$, $B$, and $c$ are defined as in i) above, i.e., given in Appendix D below.
    We report for different values of iterations $m$ and orders of graph inputs $n$, how close the two outputs of the Pfaffian GNNs are, on tree inputs $T[0,k]$ and $T[1,k]$ are. Since the root of these trees have respectively 00 and $1$ as output of the query $Q$, the ReLU GNN constructed will have a constant gap of $1$. Theorem 3.2 states that these outputs will tend to be arbitrarily close when the order of the graph increases, for a fixed $m$. In addition, Theorem 3.4 tells us that the number of iterations should impact exponentially well how large graphs can be handled to be at a given level of precision of a gap $1$.
@@
Figure 3: Separation power of $\operatorname{CReLU}$ GNN vs. step-like Pfaffian GNNs as a function of the size of the input.
@@ 6.3 Empirical results
Figure 3 reports our results for the first experiment. As our theory predicts, the $\operatorname{CReLU}$ GNN designed with the appropriate architecture and weights is able to distinguish the source vertices of both trees with constant precision 1, regardless of the number of vertices. On the other end, the Pfaffian GNNs struggle to do separation for values of $m\leq 8$ (recall that $m$ is the number of composition, i.e. the number of layers in the combinations functions) for graphs of less that less than 1500 vertices. As the value of $m$ increases slightly to $10$ for the first query and $12$ for the second, the capacity of these GNNs dramatically increases to reach a plateau close to 1 for graphs up to $4000$ vertices. This also illustrates second our second theoretical result prediciting that the step-like activation functions endow GNNs efficient approximation power of GC2 queries.
Figure 4 reports the mean square error on our test set for learning the same two queries. We observe that the error is similar for both GNNs, with a slight advantage for Pfaffian GNNs. This suggests that better uniform capacities does not imply that learning the queries is getting easier, given the same architecture and learning method. In particular, we conjecture that the set of weights such that the ReLU GNN computes the query is challenging to reach with standard optimization learning techniques based on variants of stochastic gradient descent.
@@
Figure 4: Learning queries with Pfaffian GNN vs ReLU GNN. The mean square error per graph on the test set is displayed as a function of the order of the graph.
@@ 7 Discussion and open questions
Sammy Khalife gratefully acknowledges support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341, and from the Office of Naval Research (ONR) grant N00014-24-1-2645. J.T.-C. thanks Jazz G. Suchen for useful discussions during the write-up of this paper, particularly regarding Proposition C.3.
@@ Appendix A Facts about Pfaffian Functions
The first proposition is just a direct consequence of the so-called o-minimality of Pfaffian functions Wilkie (1999), which guarantees that Pfaffian systems share with polynomial systems their finiteness properties—among which we have the finiteness of zero-dimensional zero sets. The first lemma is an easy consequence of the chain rule and the definition of Pfaffian functions. The second lemma follows easily from the proposition, but we prove it since the proof is not straightforward.
@@
The set of Pfaffian functions is stable by composition.
@@
Since $f$, $f^{\prime}$ and $f^{\prime\prime}$ are Pfaffian, by Proposition A.1, they have a finite number of zeros—if any of them are constant, then $f$ cannot be bounded. Hence there is $R>0$ such that all the zeros of $f$, $f^{\prime}$ and $f^{\prime\prime}$ are inside $(-R,R)$. In this way, since $f^{\prime}$ is bounded on any bounded interval, we only need to prove that $f^{\prime}$ is bounded on $[R,\infty)$ and on $(-\infty,-R]$. Without loss of generality, we will focus on the former case.
@@
Since
$\tanh(x)=2\operatorname{Sigmoid}(2x)-1,$ (2)
we only need to show that $\operatorname{Sigmoid}$ is Pfaffian, by Lemma A.2. Now, $\operatorname{Sigmoid}^{\prime}(x)=-e^{-x}\operatorname{Sigmoid}(x)^{2}$. Therefore
$\operatorname{Sigmoid}^{\prime}(x)=P(e^{-x},\operatorname{Sigmoid}(x)),\,e^{-x}=Q(e^{-x})$
with $P(U,V)=-UV^{2}$ and $Q(Z)=-Z$, and so $\operatorname{Sigmoid}$ is Pfaffian. ∎
@@
$H:=\{\sigma:\mathbb{R}\rightarrow\mathbb{R}:\sigma\text{ is Pfaffian, bounded and $\sigma$ converges fast in $+\infty$}\}$
Let $S_{r}$ be the set of $\mathbb{R}^{r}\rightarrow\mathbb{R}$ functions of fast decay in the last variable:
@@
where for functions $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $\lim_{z\rightarrow+\infty}F(z)=0$ means that for every $\epsilon>0$ there exists $\exists\beta\geq 0$, such that $\min(z_{1},\cdots,z_{d})\geq\beta\implies\lvert F(z)\rvert<\epsilon$.
For non-negative integers $k,m$, let $T[x,k,m]$ be the rooted tree of Figure 2. For ease of notation, if the context allows it, we will refer to $T[x,k,m]$ simply as $T$. For any non-negative integer $t$, $v\in V(T)$, let $\xi^{t}(v,T)$ be the embedding returned for node $v$ by a GNN with activation function $\sigma\in H$ after $t$ iterations. Let $M>0$ be an integer. We will prove by induction on $t$ that for any $t\in\{1,\cdots,M\}$:
@@
with the following properties: a) $v_{t}$ is constant with respect to $x$ and $m$. b) $g_{t}$ and $h_{t}$ are bounded Pfaffian functions that depend only on the combination and aggregation functions and the iteration $t$, and c) For every $t\in\{1,\cdots,M\}$, $\nu_{t},\eta_{t},\epsilon_{t},\epsilon^{1}_{t},\nu^{1}_{t}\in S_{3}$.
@@
- i) the above system of equation is verified
- ii) $g_{u}$, $h_{u}$ are Pfaffian and bounded functions
- iii) $\nu_{u},\eta_{u},\epsilon_{u},\epsilon^{1}_{u}$ and $\nu^{1}_{u}\in S_{3}$
Induction step: First note that $\mathsf{comb}$ has bounded derivatives as the derivative of a bounded Pfaffian function is also bounded. Let $K_{\mathsf{comb}}:=\sup_{x\in\mathbb{R}^{2}}\lVert\partial_{x_{1}}\mathsf{comb}(x),\partial_{x_{2}}\mathsf{comb}(x)\rVert_{2}$.
@@
$=\mathsf{comb}(h_{t}(k)+\nu_{t,m}(k),g_{t}(k)+\epsilon_{t,m}(k))$
$:=\mathsf{comb}(h_{t}(k),g_{t}(k))+\nu_{t+1,m}(k)$
$:=h_{t+1}(k)+\nu_{t+1,m}(k)$
@@
Using Corollary A.1, for every $u\in\{1,\cdots,t+1\}$, $g_{u}$ is either constant equal to 00 on $\mathbb{R}$ or has no zeroes after some rank. Therefore, there exists an integer $k^{\star}$ such that for every $k_{i}\geq k^{\star}$, for every $u\in\{1,\cdots,t+1\}$:
- Case a) $g_{u}(k_{i})>0$
- Case b) or $g_{u}(k_{i})<0$
- Case c) $g_{u}=\tilde{0}$ on $\mathbb{R}$
(i.e. we can select a common $\beta=k^{\star}$ for each iteration). Now, suppose that $\lvert k\rvert\geq\beta$.
By continuity of $\mathsf{comb}$, $g_{t}$ having a limit in +$\infty$+ ∞,
@@
where $l=\lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)$ if case a) for $u=t+1$ and $l=\lim_{y\rightarrow-\infty}\mathsf{comb}(v_{t},y)$ if case b) for $u=t+1$, and $l=\mathsf{comb}(v_{t},0)$ otherwise (case c) for $u=t+1$). Hence
@@
=$=\begin{cases}\lim_{y\rightarrow-\infty}\mathsf{comb}(v_{t},y)\;\text{if $g_{t}(k^{\star})<0$ }\\ \lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)\;\text{if $g_{t}(k^{\star})>0$ }\\ \mathsf{comb}(v_{t},0)\;\text{otherwise}\end{cases}\qquad+\eta_{t+1}(x,k,m)$
where $\eta_{t+1}=\eta^{1}_{t+1}+\eta^{2}_{t+1}$ (in the following, we suppose without loss of generality that $g_{t}(k^{\star})>0$) with
@@
$\lvert\eta^{(2)}_{t+1}(x,k,m)\rvert:=\lvert\mathsf{comb}(v_{t},g_{t}(x)+(m-1)g_{t}(k))-\lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)\rvert$
In this way,
$\lvert\eta^{1}_{t+1}(x,k,m)\leq K_{\mathsf{comb}}\sqrt{2}\lVert(\eta_{t}(x,k,m),\epsilon^{1}_{t}(x,k,m)+(m-1)\epsilon_{t}(x,k,m))\rVert_{\infty}$
The error term gets multiplied by $m$. We know by the induction hypothesis that for every polynomial $P$
$\forall\epsilon>0,\exists\beta,\;\forall m\geq\beta,\;\;\beta\leq\lvert k\rvert\implies\lvert P(m)\epsilon_{t}(x,k,m)\rvert\leq\epsilon$
@@
The triangular inequality $\lvert\eta_{t+1,m}(k)\rvert\leq\lvert\eta^{(1)}_{t+1,m}(k)\rvert+\lvert\eta^{(2)}_{t+1,m}(k)\rvert$, this gives the desired property (after splitting the $\epsilon$ in half) and ends the induction. ∎

### v2

@@ 6.2 Experimental setup
Computing and approximating queries. For the first experiment, our set of input graphs is composed of pairs of trees $(T[0,k,m],T[1,k,m])_{k\in\mathbb{N},m\in\mathbb{N}}$ presented in Section 4. These are unicolored for $Q_{1}$ and colored as follows for $Q_{2}$: the root $s$ and the $x_{i}$ are red, and the leaves are blue. We compare for $i\in[2]$, the two following GNNs:
(a) The ReLU GNN that exactly computes the query $Q_{i}$.
We use the construction of (Barceló et al., 2020), specified in the proof of Theorem 5.1 where each of the GNNs has the same combination function in each iteration. For $Q_{1}$, the ReLU GNN has $4$ iterations (depth of $Q_{1}$); and for $Q_{2}$, the ReLU GNN has $7$ iterations (depth of $Q_{2}$). The combination functions are of the form
$\mathsf{comb}_{t}(x,y)=\operatorname{ReLU}(Ax+By+C),$
where $x,y\in\mathbb{R}^{d}$, $t\in\{0,\ldots,d-1\}$ and $d=4$ for $Q_{1}$ and $d=7$ for $Q_{2}$. The parameters $A$, $B$ and $C$ are specified in Appendix D.
(b) The Pfaffian GNN with activation function: $\sigma:x\mapsto 1/2+(2/\pi)\arctan\left(x-1/2\right)$ that expresses $Q_{i}$ non uniformly.
We use the construction in the proof of Theorem 5.1, where as in the GNNs above, the combination functions are all the same. In this case, the combination functions are of the form
$\mathsf{comb}_{t}(x,y):=\sigma^{\ell}(Ax+By+c)$
where $x,y\in\mathbb{R}^{d}$, $t\in\{0,\ldots,d-1\}$, $d=4$ for $Q_{1}$ and $d=7$ for $Q_{2}$, and the parameters $A$, $B$ and $c$ are the same as the ones of the ReLU counterparts, given in Appendix D. Note that this is a NN with $\ell+2$ layers, where $\ell$ of the layers have width $d$ and one has width $2d$.
We report for different values of $\ell$ and number of vertices of the graphs inputs (which is function of $k$ and $m$), how close the two outputs of the Pfaffian GNNs, on the tree inputs $T[0,k,m]$ and $T[1,k,m]$. Since the root of these trees have respectively 00 and $1$ as output of the query $Q$, the ReLU GNN constructed will have a constant gap of $1$. Theorem 3.2 states that these outputs will tend to be arbitrarily close when the order of the graph increases, for a fixed $\ell$. In addition, Theorem 3.4 tells us that the number of iterations should impact exponentially well how large graphs can be handled to be at a given level of precision of a gap $1$.
@@
Figure 3: Separation power of $\operatorname{CReLU}$ GNN vs. Pfaffian GNN above as function of the graph size
@@ 6.3 Empirical results
Figure 3 reports our results for the first experiment. As our theory predicts, the $\operatorname{CReLU}$ GNN designed with the appropriate architecture and weights is able to distinguish the source vertices of both trees with constant precision 1, regardless of the number of vertices. On the other end, the Pfaffian GNNs struggle to do separation for values of $\ell\leq 8$ (recall that $\ell$ is the number of compositions, i.e. the number of layers in the combinations functions) for graphs of than 1500 vertices. As the value of $\ell$ increases slightly to $10$ for the first query and $12$ for the second, the capacity of these GNNs dramatically increases to reach a plateau close to 1 for graphs up to $4000$ vertices. This also illustrates our second theoretical result prediciting that the step-like activation functions endow GNNs efficient approximation power of GC2 queries.
Figure 4 reports the mean square error on our test set for learning the same two queries. We observe that the error is similar for both GNNs, with a slight advantage for Pfaffian GNNs. This suggests that better uniform capacities does not imply that learning the queries is getting easier, given the same architecture and learning method. One potential reason is numerical: the combination functions using step-like activation functions used in our approximation of GC2 queries have likely to have large gradients. We suspect this adds to the difficulty to efficiently train these networks to optimality, despite their ability to approximate queries to an arbitrary level of precision.
@@
Figure 4: Learning queries with Pfaffian GNN vs ReLU GNN.
The mean square error per graph on the test set is displayed as a function of the order of the graph.
@@ 7 Discussion and open questions
Sammy Khalife gratefully acknowledges support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341, and from the Office of Naval Research (ONR) grant N00014-24-1-2645. Josué Tonelli-Cueto thanks Jazz G. Suchen for useful discussions during the write-up of this paper, particularly regarding Proposition C.3, and Brittany Shannahan and Lewie-Napoleon III for emotional support.
@@ Appendix A Facts about Pfaffian Functions
The first proposition is just a direct consequence of the so-called o-minimality of Pfaffian functions Wilkie (1999), which guarantees that Pfaffian systems share with polynomial systems their finiteness properties—among which we have the finiteness of zero-dimensional zero sets. The first lemma is an easy consequence of the chain rule and the definition of Pfaffian functions. The second and third lemmas follow from the proposition, but we include a proof for completeness.
@@
The addition, multiplication, composition and derivatives of Pfaffian functions are Pfaffian.
Lemma A.3.
Let $f:\mathbb{R}\rightarrow\mathbb{R}$ be a Pfaffian and bounded function. Then $f$ has a finite limit in +$\infty$+ ∞.
Proof.
By Lemma A.2, since $f$ is Pfaffian, so is $f^{\prime}$. Hence, by Proposition A.1, either $f^{\prime}$ is constant, and so $f^{\prime}=0$, or $f^{\prime}$ has finitely many zeros. In the first case, $f$ is constant and so it trivially has a limit. In the second case, let $R>0$ be such that $(R,\infty)$ does not contain any zero of $f^{\prime}$. Then either $f^{\prime}$ is strictly positive or strictly negative on $(R,\infty)$, and so either $f$ is strictly increasing or strictly decreasing on $(R,\infty)$. Hence, by the supremum axiom of the real numbers, we have that
$\lim_{x\to\infty}f(x)=\sup\{f(y)\mid y\in(R,\infty)\}$
if $f$ is strictly increasing on $(R,\infty)$, or
$\lim_{x\to\infty}f(x)=\inf\{f(y)\mid y\in(R,\infty)\}$
if $f$ is strictly decreasing on $(R,\infty)$. Since $f$ is bounded, the right-hand side is finite in both cases. Thus $f$ has a finite limit in +$\infty$+ ∞. ∎
@@
By Lemma A.2, $f^{\prime}$ and $f^{\prime\prime}$ are Pfaffian, because $f$ is so. Thus, by Proposition A.1, $f$, $f^{\prime}$ and $f^{\prime\prime}$ have a finite number of zeros—if any of them are constant, then $f$ cannot be bounded. Hence there is $R>0$ such that all the zeros of $f$, $f^{\prime}$ and $f^{\prime\prime}$ are inside $(-R,R)$. In this way, since $f^{\prime}$ is bounded on any bounded interval, we only need to prove that $f^{\prime}$ is bounded on $[R,\infty)$ and on $(-\infty,-R]$. Without loss of generality, we will focus on the former case.
@@
The only part requiring proof is that they are Pfaffian, that they are bounded is easily shown.
We have that $\tanh^{\prime}(x)=1-\tanh(x)^{2}$ and that $\operatorname{Sigmoid}^{\prime}(x)=-e^{-x}\operatorname{Sigmoid}(x)^{2}$. In this way,
$\tanh^{\prime}(x)=P(x,\tanh(x))$
with $P(x,y)=1-y^{2}$, and
$\operatorname{Sigmoid}^{\prime}(x)=P(x,e^{-x},\operatorname{Sigmoid}(x)),\,e^{-x}=Q(x,e^{-x})$
with $P(x,y_{1},y_{2})=-y_{1}y_{2}^{2}$ and $Q(x,y)=-y$, show that $\tanh$ and $\operatorname{Sigmoid}$ are Pfaffian.
Alternatively, since
$\tanh(x)=2\operatorname{Sigmoid}(2x)-1~{}\text{ and }~{}\operatorname{Sigmoid}(x)=\tanh(x/2)/2,$ (3)
it would have been enough to show that either $\tanh$ or $\operatorname{Sigmoid}$ is Pfaffian to conclude that the other is so, by Lemma A.2. ∎
@@
where the notions above where given in Definitions 2.5 and 3.2. Let $S_{r}$ be the set of $\mathbb{R}^{r}\rightarrow\mathbb{R}$ functions of fast decay in the last variable:
@@
where for functions $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $\lim_{z\rightarrow+\infty}F(z)=0$ means that for every $\epsilon>0$ there exists $\beta\geq 0$, such that $\min(z_{1},\cdots,z_{d})\geq\beta$ implies $\lvert F(z)\rvert<\epsilon$.
For non-negative integers $k$ and $m$, let $T[x,k,m]$ be the rooted tree of Figure 2 in which the root $s$ has $m$ descendants $x_{i}$, having $x_{i}$ $x$ descendants if $i=1$ and $k$ descendants if $i\geq 2$. For ease of notation, we will refer to $T[x,k,m]$ simply as $T$ and to any of the descendants of $x_{i}$ as $l_{i}$. For any non-negative integer $t$ and $v\in V(T)$, let $\xi^{t}(v,T)$ be the embedding returned for node $v$ by a GNN with activation function $\sigma\in\Xi$ after $t$ iterations. Let $M>0$ be an integer. We will prove by induction on $t$ that for any $t\in\{1,\cdots,M\}$:
@@
with the following properties: a) $v_{t}\in\mathbb{R}^{d_{t}}$ is constant with respect to $x$ and $m$. b) each coordinate of $g_{t}$ and $h_{t}$ are bounded Pfaffian functions of $k$ that depend only on the combination and aggregation functions and the iteration $t$, and c) each coordinate of $\nu_{t},\eta_{t},\epsilon_{t},\epsilon^{1}_{t},\nu^{1}_{t}$ are in $S_{3}$.
In the remaining of the proof we explicitly treat the case of one-dimensional embeddings (i.e. $d_{t}=1$ for every $t\in[T]$) for ease of notation. Our proof extends to multi-dimensional embeddings with the three additional ingredients:
i) considering only one input variable and the rest fixed, each coordinate of a combine function given by a neural network with an activation function $\sigma\in\Xi$ is a bounded Pfaffian function that is constant or superpolynomial (see Lemma A.2). Therefore, one can also use in that case Corollary A.1 on the zeroes of these functions, as we do in the one-dimensional case.
ii) the error terms of each signal can be controlled using norms of the Jacobian matrix instead of the gradient, as we shall see next.
iii) Given a bounded multivariate Pfaffian function $f$, for any fixed reals $\lambda_{1},\cdots,\lambda_{r}$, $f(\lambda_{1}m,\lambda_{2}m,\cdots,\lambda_{r}m)$ has a limit when $m\rightarrow+\infty$. This follows from the fact that the univariate function $x\mapsto f(\lambda_{1}x,\lambda_{2}x,\cdots,\lambda_{r}x)$ is Pfaffian bounded.
@@
  1. i)
    The above system of equation is verified.
  2. ii)
    $g_{u}$, $h_{u}$ are Pfaffian and bounded functions.
  3. iii)
    $\nu_{u},\eta_{u},\epsilon_{u},\epsilon^{1}_{u}$ and $\nu^{1}_{u}\in S_{3}$.
Induction step: For the remaining of the proof, for convenience we drop the dependency on $t$ of $\mathsf{comb}_{t}$ and simply write $\mathsf{comb}$, which is a neural network with some activation $\sigma\in\Xi$. First note that $\mathsf{comb}$ has bounded derivatives as the derivative of a bounded Pfaffian function is also bounded, see Lemma A.4. Let $K_{\mathsf{comb}}:=\sup_{x\in\mathbb{R}^{2}}\lVert\partial_{x_{1}}\mathsf{comb}(x),\partial_{x_{2}}\mathsf{comb}(x)\rVert_{2}$.
@@
$=\mathsf{comb}(h_{t}(k)+\nu_{t}(x,k,m),g_{t}(k)+\epsilon_{t}(x,k,m))$
$:=\mathsf{comb}(h_{t}(k),g_{t}(k))+\nu_{t+1}(x,k,m)$
$:=h_{t+1}(k)+\nu_{t+1}(x,k,m)$
@@
Using Corollary A.1, for every $u\in\{1,\cdots,t+1\}$, $g_{u}$ is either constant equal to 00 on $\mathbb{R}$ or has no zeroes after some rank. Therefore, there exists an integer $k^{\star}$ such that for every $k_{i}\geq k^{\star}$ and every $u\in\{1,\cdots,t+1\}$, one of the following cases holds:
  1. Case a)
    $g_{u}(k_{i})>0$.
  2. Case b)
    $g_{u}(k_{i})<0$.
  3. Case c)
    $g_{u}=\tilde{0}$ on $\mathbb{R}$.
We remind the reader that a similar case discussion here extends to the multidimensional case, i.e. when the function $g_{u}$ take vectorial values, and when the combine function has multi-dimensional co-domain, see for example comment iii). Note that we can select a common $\beta=k^{\star}$ for each iteration. Now, suppose that $\lvert k\rvert\geq\beta$.
By continuity of $\mathsf{comb}$ and $g_{t}$ having a limit in +$\infty$+ ∞,
@@ 3. Case c)
where $l=\lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)$ if case a) and $l=\lim_{y\rightarrow-\infty}\mathsf{comb}(v_{t},y)$ if case b), and $l=\mathsf{comb}(v_{t},0)$ otherwise (case c). Hence
@@
=$=\begin{cases}\lim_{y\rightarrow-\infty}\mathsf{comb}(v_{t},y)+\eta_{t+1}(x,k,m)\;&\text{if $g_{t}(k^{\star})<0$ }\\ \lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)+\eta_{t+1}(x,k,m)\;&\text{if $g_{t}(k^{\star})>0$ }\\ \mathsf{comb}(v_{t},0)+\eta_{t+1}(x,k,m)\;&\text{otherwise}\end{cases}$
where $\eta_{t+1}=\eta^{1}_{t+1}+\eta^{2}_{t+1}$. In the third case the induction step goes through immediately. In the following, we suppose without loss of generality that $g_{t}(k^{\star})>0$, with
@@
$\lvert\eta^{(2)}_{t+1}(x,k,m)\rvert:=\left\lvert\mathsf{comb}(v_{t},g_{t}(x)+(m-1)g_{t}(k))-\lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)\right\rvert$
First, it is clear that $\eta^{2}_{t+1}(x,k,m)$ can be made less than $\epsilon$ for some $m$, due to the existence of the limit $\lim_{y\rightarrow+\infty}\mathsf{comb}(v_{t},y)$ and that $g_{t}(k)$ is chosen to be non zero (for every $k$ after rank $k^{\star}$ chosen above). Furthermore, $\eta^{(2)}_{t+1}$ is still in $S_{3}$ as for every fixed $x,k\geq k^{\star}$ the function $m\mapsto\mathsf{comb}(v_{t},g_{t}(x)+(m-1)g_{t}(k))$ tends superpolynomially fast to its limit, due to the assumption made on the activation function and again from the fact that $g_{t}(k)\neq 0$ for every $k\geq k^{\star}$. Second,
$\lvert\eta^{1}_{t+1}(x,k,m)\rvert\leq K_{\mathsf{comb}}\sqrt{2}\lVert(\eta_{t}(x,k,m),\epsilon^{1}_{t}(x,k,m)+(m-1)\epsilon_{t}(x,k,m))\rVert_{\infty}$
We see that for this part the error terms $\epsilon^{1}_{t}$ and $\epsilon_{t}$ get multiplied by $m$. We know by the induction hypothesis that for every polynomial $P$,
$\forall\epsilon>0,\exists\beta,\;\forall m\geq\beta,\;\;\beta\leq\lvert k\rvert\implies\lvert P(m)\epsilon_{t}(x,k,m)\rvert\leq\epsilon.$
@@
By the triangle inequality, $\lvert\eta_{t+1,m}(k)\rvert\leq\lvert\eta^{(1)}_{t+1,m}(k)\rvert+\lvert\eta^{(2)}_{t+1,m}(k)\rvert$, and so the desired property follows (after splitting the $\epsilon$ in half) and ends the induction. ∎
Remark B.1.
[…]